OpenStudy (anonymous):

Please Help! MEDAL 1. Complete the table below to solve the equation 2.5x − 10.5 = 64(0.5x). x f(x) = 2.5x − 10.5 g(x) = 64(0.5x) 2 3 4 5 6

OpenStudy (trevor12):

2.5x–10.5=64(0.5x) Distribute 64: 2.5x–10.5=32x Subtract 2.5x to both sides: −10.5=29.5x Divide 29.5 to both sides: x=−0.355932203